OpenStudy (anonymous):

What is the y-intercept for this linear equation? y = 5x - 10

OpenStudy (anonymous):

It's in the form y = mx+, where m= gradient and c= y-intercept

OpenStudy (anonymous):

So the gradient/slope would be five and the y-intercept would be -10

OpenStudy (anonymous):

but i need to showe work

OpenStudy (anonymous):

That is all the work you'd need to show

OpenStudy (luigi0210):

y intercept is when x=0, so plug zero in and solve

OpenStudy (luigi0210):

y=5x-10 y=5(0)-10 y=-10

OpenStudy (anonymous):

but for this one is y-intercept

OpenStudy (anonymous):

Does it ask you the y-intercept only, or are you required to graph it?

OpenStudy (anonymous):

i am ganna posted more